Circular No. 7352

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000C IN NGC 2415 E. Cappellaro and M. Turatto, Osservatorio Astronomico di Padova,

report that spectra of SN 2000C (cf. IAUC 7348) have been obtained by A. Pizzella, E. M. Corsini, and D. Moro with the Asiago 1.22-m telescope (+ Boller & Chivens spectrograph) on Jan. 25.9 UT and by G. Galletta with the 1.8-m telescope (+ AFOSC) on Jan. 27.9 (range 350–750 nm, resolution 2 nm). The calibrated spectra are very similar to that of the type-Ic supernovae 1987M and 1990aa, 1–2 weeks after maximum (cf. Filippenko et al. 1992, Ap.J. 384, L37), with prominent Ca II and Fe II features. Also visible is an absorption at 625 nm that, if attributed to Si II 635.5-nm (as in the case of SN 1987M), indicates an expansion velocity of ∼ 9000 km/s. Therefore, despite the apparent coincident location, SN 2000C is a distinct object from SN 1998Y.

S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2000C, taken by P. Berlind on Jan. 28.14 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-Ic supernova near maximum light. The spectrum shows a blue continuum with broad absorption features of Ca II, Mg II, Fe II, Na I, Si II, and O I, and is similar to maximum-light spectra of SN 1987M and SN 1994I (Filippenko 1997, Ann. Rev. A.Ap. 35, 309). Superimposed narrow emission lines yield a host galaxy recession velocity of 3710 km/s.

SUPERNOVA 2000D IN UGC 1767 Jha et al. also report that a spectrum of SN 2000D, taken by Berlind on

Jan 28.09 UT (as above), shows this to be a type-II supernova at an early epoch. The spectrum exhibits a very blue continuum with P-Cyg Balmer lines and He I. Superimposed Hα emission yields a recession velocity for the host galaxy of 5280 km/s. The supernova expansion velocity, measured to the Hβ absorption trough, is 11 400 km/s.

1E 161348–5055 Corrigenda to IAUC 7350. On line 15, for Aug. 18 read Aug. 18.11.

On the last line, for compact supernova remnant.” read compact remnant of a supernova explosion.”

2000 January 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green